1. If the nominal gross domestic product (GDP) for the year 2000 was $6.2 trillion and...

1.

If the nominal gross domestic product (GDP) for the year 2000 was $6.2 trillion and the price index was 200, the real gross domestic product (GDP) for 2000 was _____.​

$12.4 trillion​

$6.2 trillion​

​$3.1 trillion-

$24.3 trillion​

$18.6 trillion​

2.Which of the following is not a problem for less-developed countries?

Group of answer choices

Low labor productivity.

Poor health and nutrition.

Shortages of labor.

High unemployment rates.

low life expectancy

3.The economy is fully employed when there is no:

Group of answer choices

seasonal unemployment.

cyclical unemployment.

structural unemployment.

frictional unemployment.

4.Use the aggregate expenditures model and assume an economy is in equilibrium at $5 trillion which is $250 billion below full-employment GDP. If the marginal propensity to consume (MPC) is 0.60, full-employment GDP can be reached if government spending:

Group of answer choices

increases by $250 billion.

increases by $100 billion.

is held constant.

inreases by $60 billion.

5.Which of the following can be classified as a regressive tax?

Group of answer choices

Sales tax.

All of these.

Gasoline tax.

Excise tax.

6.Whom does the Bureau of Labor Statistics include in the "discouraged workers" group?

Group of answer choices

Members of the noninstitutional population who say they want to be employed but aren't searching for a job.

All those between 16 and 65 who are neither disabled nor in an institution and are also neither employed nor seeking employment.

Civilian labor force members who have chosen early retirement because they dislike their work or think the pay is too low.

Members of the civilian labor force who are looking for a job but cannot find one.

Anyone in the population who is not working.
Answer #1

1) Real GDP=( Nominal GDP/Price Index)×100

Real GDP= (6.2/200)×100= $3.1 trillion

Hence, third option is correct.

2) shortages of labour is not a problem for less developed countries as there is a plenty of labour and shortage of capital.

Hence, third option is correct

3) The economy is fully employed when there is cyclical unemployment(employment caused by business cycle).

Hence, second option is correct.

4) Spending Multiplier= 1/(1–MPC)= 1/(1–0.60)= 2.5

Required Increase in Government Spending= Required Increase in GDP/Spending Multiplier= 250/2.5= $100 billion

Hence, full employment GDP can be reached if government spending increased by $100 billion.

Therefore, second option is correct.

5) all the given kind of taxes, that is, sales tax, gasoline tax and excise tax are regressive.

Hence, second option is correct.

6) Bureau of labour statistics include in the discouraged workers groups all those between 16 and 65 who are neither disabled nor in an institution and are also neither employed Nor seeking employment.

Hence, second option is correct